OpenStudy (anonymous):

y^2 -49 y-7 ------- div ---- y y+9 divide and simplify

OpenStudy (anonymous):

(y+7)(y+9) -------- y

OpenStudy (anonymous):

quotient y+16 remainder 63

OpenStudy (anonymous):

y^2-49 y-7 -------- div ----- y y+9 (y+7)(y-7) y+9 = -------- * ----- y y-7 (y+7)(y+9) =------ ----- y y^2+16y+63 =------------- y